OpenStudy (anonymous):

How do I use partial fractions to solve for tan(x/2)/((x+2)(x+3))?

OpenStudy (anonymous):

Is it possible? I thought partial expansion is only for polynomials

OpenStudy (anonymous):

I want to integrate this function with respect to dx

OpenStudy (anonymous):

you can not do partial fractions. you COULD use convolutions though, although i don't know what class you're currently in

OpenStudy (anonymous):

I'm actually attempting complex values..

OpenStudy (anonymous):

where x is actually z but I don't know if you guys could help me.. Yeah I know convolution but that's not the class I learnt it in

OpenStudy (anonymous):

what class is this?

OpenStudy (anonymous):

i think you might have done something wrong earlier if you have to take that integral... it would get pretty nasty to do so

OpenStudy (anonymous):

Oh I see what I did wrong.. okay it's actually integral of tan(z/2) / (z^4-16) dz

OpenStudy (anonymous):

it's complex math analysis, do you have any ideal on how to integrate that?

OpenStudy (anonymous):

no, and wolfram alpha does not know how to either

OpenStudy (anonymous):

it says use partial fractions! that's why I'm so confused

OpenStudy (anonymous):

oh okay thanks for your help anyways

OpenStudy (anonymous):

you must be doing something wrong earlier, because you cannot use partial fraction decomposition

OpenStudy (anonymous):

not that i know how to at least

OpenStudy (anonymous):

no we were given tan ..

OpenStudy (anonymous):

it's alright I'm sure theres a way to reduce it I just haven't seen it yet

OpenStudy (anonymous):

Thanks for your help though

OpenStudy (primeralph):

Still need help?

OpenStudy (anonymous):

You could write \[\frac{\tan\frac{x}{2}}{(x+2)(x+3)}=\tan\frac{x}{2}\left(\frac{A}{x+2}+\frac{B}{x+3}\right)\] I'm not sure if that would get you anywhere, though.

OpenStudy (anonymous):

I get it now thanks for all your help.

OpenStudy (anonymous):

Yes that's how you go about it sithsandgiggles,